Two-point charges are arranged on the x-y coordinate system as follows: q1= 4x10^-8 C at (0,3 m) and q2= -9x10^-7 C at (5,0 m). Find the electric origin at (0,0). Using the Pythagorean theorem calculate for the magnitude E of the resultant electric field.


1
Expert's answer
2022-02-16T08:36:06-0500

Answer

Electric field at origin due to q1

"E_1=\frac{kq_1}{r^2}\\=\frac{9*10^9*4*10^{-8}}{3^2}\\=40N/C"

Electric field at origin due to q2

"E_2=\frac{kq_2}{r^2}\\=\frac{9*10^9*9*10^{-7}}{5^2}\\=324N/C"

So net electric field at origin is "E=\sqrt{E_1^2+E_2^2}\\=\sqrt{(40)^2+(324)^2}\\=326.46N/C"